OpenStudy (anonymous):

Can someone please help me find the point on the line 2x+3y+4=0 which is closest to the point (3,-2).

OpenStudy (anonymous):

Ok This is going to sound weird but when you write any point and a line in a piece of paper and try to find the point of the line where it's closest to that point what happens?

OpenStudy (anonymous):

Draw a line between the two points.

OpenStudy (anonymous):

a little confused

OpenStudy (anonymous):

do you use the distance forula and the point given then somehow get another point plug in and take derivative to maximize?

OpenStudy (anonymous):

\[2x+3y+4=0 \]closest to \((3,-2)\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok first lets solve for y in the line to get \(y=-\frac{2}{3}x-\frac{4}{3}\) so any point on the line will look like \((x, -\frac{2}{3}x-\frac{4}{3})\)

OpenStudy (anonymous):

then distance squared between a point on the line and \((3,-2)\) will be \[(x-3)^2+(-\frac{2}{3}x-\frac{4}{3}+2)^2\]

OpenStudy (anonymous):

thats as far as i got and then i got stuck

OpenStudy (anonymous):

ok we have \[(x-3)^2+(-\frac{1}{2}x-\frac{2}{3})^2\] now square and combine like terms first

OpenStudy (anonymous):

my god is this ugly. where did this problem come from? maybe there is an easier way, but once we have come this far might as well finish

OpenStudy (anonymous):

yea its definitely not an eye pleasing problem in terms of integers

OpenStudy (anonymous):

then we can do it the easier way. looks like we get \[\frac{5}{4}x^2-\frac{16}{3}x+\frac{85}{9}\] http://www.wolframalpha.com/input/?i=%28x-3%29^2%2B%28-x%2F2-2%2F3%29^2

OpenStudy (anonymous):

derivative is \[\frac{5}{2}x-\frac{16}{3}\] set equal zero and solve for x

OpenStudy (anonymous):

so quadratic formula im guessing

OpenStudy (anonymous):

i get \[x=\frac{32}{15}\]

OpenStudy (anonymous):

no not quadratic formula, it is a line

OpenStudy (anonymous):

kk let me check

OpenStudy (anonymous):

I got 31/13 doing something completely different :P

OpenStudy (anonymous):

wolfram gives the same answer

OpenStudy (anonymous):

oh damn damn damn i made a mistake hold on!

OpenStudy (anonymous):

ok its correct 31/13 , 2.923 are the cordinates thanks guys

OpenStudy (anonymous):

i wrote it in wrong, sorry. let me try again square of the distance is \[\frac{13}{9}x^2-\frac{46}{9}x+\frac{85}{9}\]

OpenStudy (anonymous):

sorry -2.923

OpenStudy (anonymous):

derivative is \[\frac{26}{9}x-\frac{46}{9}\] set = 0 and solve get \[x=\frac{46}{26}=\frac{23}{13}\] did i make another mistake?

OpenStudy (anonymous):

yes i see it

OpenStudy (anonymous):

I'm not sure but what I did was make a function using point slope. I used the point that was given and the inverse slope of the function given. Used that function set it equal to the given function and found the point in where both cross each other.

OpenStudy (anonymous):

should be \[(x-3)^2+(-\frac{2}{3}x+\frac{2}{3})^2\]

OpenStudy (anonymous):

@Romero yes that is the best method. slope will be perpendicular

OpenStudy (anonymous):

There is always one way to solve it!! :)

OpenStudy (anonymous):

slope of this line is \[-\frac{2}{3}\] slope of perpendicular line is \[\frac{3}{2}\] equaton of the line is \[y+2=\frac{3}{2}(x-3)\] or \[y=\frac{3}{2}x-\frac{5}{2}\] and then find the point of intersection